OpenStudy (anonymous):

Is this correct? y=-2(x+1)^2+4 - reflects x axis, shifts 1 left, up 4, and has a verticle compression??

OpenStudy (zpupster):

looks right

OpenStudy (amistre64):

y = (x)^2 y = -(x)^2 , reflects y = -2(x)^2, this makes things longer in the vertical so im leary of the vertical "compression"; but then im out of date with the terminology y = -2(x+1)^2 , left by 1 y = -2(x+1)^2+4 , up by 4

OpenStudy (amistre64):



OpenStudy (amistre64):

vertical stretch does seem more appropriate http://regentsprep.org/Regents/math/algtrig/ATP9/funclesson1.htm

OpenStudy (anonymous):

thank you so much for helping!

OpenStudy (anonymous):

that's where I get confused

OpenStudy (amistre64):

when we multiply by some constant whose absolute value is greater than 1 |-2| > 1 we are moving the points further away from the x axis. y=x; when x=1, y=1 y=2x; when x=1, y=2 y=2 is further away from the x axis then y=1

OpenStudy (amistre64):

when we mutiply by a constant that is a proper fraction; we are essentially dividing ... and lessening the distance

OpenStudy (amistre64):

y=x, when x=1, y=1 y=1/2 x, when x=1, y=1/2 notice that y=1/2 is closer to the x axis than y=1

OpenStudy (amistre64):

compression is moving things closer, stretching moves them further away ..

OpenStudy (anonymous):

so... than this problem is compression, no? .__.

OpenStudy (amistre64):

compare it: y = x^2 y = -2x^2 for any value of x (other than 0 in this case), how does that effect the y value?

OpenStudy (anonymous):

.... D: idk :(

OpenStudy (anonymous):

the closer it is to 1, the skinnier the graph gets?

OpenStudy (amistre64):